[clas_members] [New Logentry] RC Daily Summary August 15, 2022

yprok at jlab.org yprok at jlab.org
Mon Aug 15 10:50:01 EDT 2022

Logentry Text:
Sunday, August 14

Day Shift: Took data with NH3(+).  Changed HWP to OUT at 1pm. Target polarization ~51%. Bleedthru test showed ~1nA current from Hall C.

Swing  and Owl Shifts: Took data with NH3(+), HWP OUT

Monday developments:  Hall C will decide when they will do pass change today at noon, so it is still not certain at this moment.  
PD informed us that the current plan, assuming RadCon approves, is to turn on the vertical separator and eliminate the bleedthrough that way this afternoon once Halls C and D are finished. At the moment Hall D has beam and Hall C is down. 

Cannot make any definitive plans at the moment; still waiting on several timelines. Informed the the target group that we are expecting downtime sometime this afternoon so they could do an anneal. 

We will have the daily RC meeting today at 2:00pm

RC Meeting Coordinates:
Meeting ID: 160 867 4253
Passcode: 786506

Data taking progress:
Day 4.0 nA on NH3    completed runs:    HWP IN :  16682(67M)    HWP OUT: 16683(68M)
Swing 4.0 nA on NH3 (HWP OUT): 16685(12M),16686(37M), 16688(72M
Owl 4.0 nA on NH3 (HWP OUT):     16689(17M), 16690(7M), 16692(57M)


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4029193
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220815/eb30168c/attachment-0001.html>

More information about the clas_members mailing list